OpenStudy (anonymous):

The ratio of the perimeters of two rectangles is 4 to 9. The perimeter of the larger rectangle is 54 inches. What is the perimeter of the smaller rectangle? A. 121.5 inches B. 13.5 inches C. 36 inches D. 24 inches

OpenStudy (anonymous):

i think D

OpenStudy (anonymous):

A lot of the questions you're asking have the same concept. You really should learn it so you can answer these without the internet. 4/9=x/54 x=24

OpenStudy (johnweldon1993):

You think correctly!

OpenStudy (whpalmer4):

if the ratio of small to large is 4:9, then \[\frac{\text{small}}{54} = \frac{4}{9}\]

OpenStudy (anonymous):

thanks